Te ttal number of ornge football helmets ad blue football helmts is 44. If twice the numbe of blue fotball helmets exceeds the number of orange football helmets by 4, how many are there of each color?
Answer by bam878s(77)

Let x = # of orange helmets and y = number of blue helmets.
Since there are 44 total we know that x + y = 44.
Now, we also know that twice the number of blue helmets exceeds the number of orange helmets by 4. We can write this as:
2y - 4 = x.
Now we can plug this value for x back into the original equation and solve for y
(2y-4) + y = 44
3y - 4 = 44
3y = 48
y = 16. Now we can use this value to solve for x
x + 16 = 44
x = 44 - 16
x = 28.
So x = 28 and y = 16 and thus, there are 28 oragne helmets and 16 blue helmets.
Hope this helps!!